OpenStudy (anonymous):

What happens to the volume of a cube if the length of an edge is doubled? options: It is multiplied by 2. It is multiplied by 4. It is multiplied by 6. It is multiplied by 8.

OpenStudy (anonymous):

2^3

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

no 2^3 = 8

OpenStudy (anonymous):

what about this one?

OpenStudy (anonymous):

I can't answer because I'm masturbating :DDD

OpenStudy (michele_laino):

for example is the side of your cube is a, then the volume is a^3. If the side of your cube is 2*a, then the volume of your cube is (2*a)^3=8*a^3

OpenStudy (anonymous):

wtf seriously just ask for nudes like everyone els @sourwing

OpenStudy (anonymous):

Thx @Michele_Laino

OpenStudy (michele_laino):

Thank you!